OpenStudy (anonymous):

Solve by substitution: 2x+5y=-4 3x-y=11

OpenStudy (gw2011):

Let 2x=-5y-4 x=-(5/2)y-2 3x-y=11 3[-(5/2)y-2]-y=11 -(15/2)y-6-y=11 -(17/2)y-6=11 -(17/2)y=11+6 -(17/2)y=17 y=(17)[-(2/17)] y=-2 Substitute this into the following: 2x+5y=-4 2x+5(-2)=-4 2x-10=-4 2x=-4+10 2x=6 x=3